### v1

@@ Abstract
Collective estimation manifests computational intelligence emerging from inter-individual local interactions, e.g., by aggregating opinions from neighbors to estimate a quantity. Use cases of collective estimation may include directed motion in physical space, such that agents, for example, have to collectively explore a distributed feature, and collectively agree on a numerical value. In doing so, collectives face several challenges in achieving precise estimations. These challenges exhibit complex behaviors, particularly when the interaction network and opinion of agents evolve simultaneously. We take homophilic networks as an example, where disproportionate interaction with like-minded neighbors leads to the emergence of echo chambers, preventing collective consensus. Our simulation results confirm that, besides a lack of exposure to attitude-challenging opinions, seeking reaffirming information entraps agents in echo chambers. We propose a generic novel approach based on a Dichotomous Markov Process (DMP) where stubborn agents (called Messengers) connect the disconnected clusters by physically transporting their opinions to other clusters to inform and direct the other agents. We show that diverse collective behaviors arise from the DMP and study a continuum between task specialization with no switching (full-time Messengers), generalization with slow task switching (part-time Messengers), and rapid task switching (short-time Messengers) and its impact on system performance. Our results show that stubborn agents can, in various ways, break the echo chambers and promote consensus in collective opinion.
@@ 1 Introduction
Collective behaviors exhibited by animal collectives or groups of interacting artificial agents are fascinating examples of self-organization. They enable groups to solve problems collectively, that cannot be solved by any individual alone. This phenomenon is often referred to as collective or swarm intelligence [1, 2, 3]. While significant progress has been made over the past decades in understanding the fundamentals of collective intelligence, many questions remain open regarding actual mechanisms underlying collectively intelligent behavior, in particular in spatially embedded systems lacking the capability for global information exchange between individual agents.
Among the many different exhibitions of intelligence in collectives, the wisdom of crowds effect [4, 5, 6] stands out as a great candidate for studying collective computational intelligence [3]. While the core idea–that the average of many imperfect estimations can be remarkably close to the true value–seems straightforward, achieving wise and precise estimations depends on meeting certain criteria [5, 7, 8, 9, 10]. Typically, it assumes full knowledge and all estimations are fully accessible, which is not naturally feasible in decentralized settings. However, distributed consensus models, particularly DeGroot-like models [11], implement the wisdom of crowds effect without a centralization assumption [12, 7]. These models also provide mechanisms for how opinions spread across the interaction network, leading to behaviors known as opinion dynamics [13, 14]. Achieving a consensus has shown potential applications in collective estimation scenarios and can be used in a variety of engineering fields ranging from machine learning to networks of sensors and groups of robots [15, 16, 17]. Moreover, studying these behaviors gives insights into how similar complex systems work, as used in economics, politics, and social sciences [18, 14, 7, 19, 20]. Modeling behavior on the individual level helps in explaining the dynamics of opinion spread in collectives. These models usually follow a very simple set of rules; Examples are Voter models, bounded confidence, and gossip methods [21, 22, 23]. The underlying network of interaction among agents is an important determinant of collective behavior [23, 24]. For example, we know that the connectivity of the network controls the tradeoff between speed and accuracy of consensus formation in collectives with static interaction networks–stronger connected networks promote faster consensus decisions at the expense of decision accuracy [17, 10].
@@ 1.1 Homophily and Echo Chambers
Contrary to the assumption of static networks in early models of opinion dynamics, many systems display dynamic networks, as agents tend to constantly rearrange their connections with their peers, making the collective behavior more complex [25]. These rearrangements can be described following different rules specifically defined for the network dynamics; for instance, consider homophily: the disproportional tendency of agents to establish links with like-minded neighbors [26, 27, 28]. By adding a new dynamic to the interactions, and hence altering how the information flows in the network, homophily changes the resulting collective behavior. One consequence of homophily in network dynamics is the emergence of “echo chambers”: clusters that are highly homogeneous internally and very heterogeneous compared to each other. The strong effects of homophily and echo chambers on collectives highlight their critical importance; particularly, when these effects lead to the emergence of phenomena that are considered threats to human society [29]. For instance, their effect on the spread of misinformation [30, 31], where the homogeneity of information within clusters acts as a driver to the diffusion of misinformation [29, 32, 33]. Ultimately, echo chambers reinforce only the existing perspective, foster confirmation biases, and prevent clusters from receiving attitude-challenging information [32, 34, 30]. Other consequences of homophily include the formation of filter bubbles and network segregation [19, 33]. Despite their adverse effects at the collective level, these echo chambers function for a purpose on the individual level: comforting agents with reaffirmation and protecting them from disagreement [32]. Although the actual adversarial outcome of echo chambers in collective dynamics is arguable [35], they indeed inhibit the flow of information on the network in our scenario where achieving consensus is an objective.
1.2 Scenario: Embedded in Space and Online Sampling of Information
@@
We introduce a novel scenario that pushes towards a physical embedding of opinion dynamics as it would be the reality for a group of mobile agents. We updated our opinion dynamic model to embed information in space. Many models have already been developed to describe a mechanistic process of how the co-evolution of homophilic networks and opinion dynamics lead to the emergence of echo chambers and complex collective behavior [36, 37, 38]. Although the opinion updating models used parallel to the network dynamics vary greatly in the literature, they presume that the initial opinions of agents come from a random pool of possible (discrete) options. Hence, these models did not focus on the origin of the opinions, but rather their evolution through social interactions. Furthermore, this elimination of information sources is inevitably reflected in the opinion-updating model, as its influence on the evolution of opinion is not further considered. Contrarily, we assume that the information agents receive comes from a specific source, and the agent’s position determines the information available to them, hence shaping their opinion. We already know that the position of agents in the network matters [39], but so does the position with respect to the information source. Therefore, we add the information space in our model to account for its effects on shaping opinions. This space resembles the source of information available to agents and is modeled as a 2D Euclidian space, similar to our previous work [17]. Compared to the previous discrete models of initial opinions in homophilic networks, the information landscape provides continuous opinions with arbitrary distributions throughout simulations.
@@ 1.3 A Motion Model: Homophily as a Pulling Force
Given our agents are embedded in an information space, we require new rules to implement homophily in space. Similarly, we define neighborhood as the local proximity (determined by communication range) of agents in the information landscape. This definition of communication range compares to the thresholds in bounded confidence models of opinion dynamics [22, 40]. In the original models of homophily without spatial constraints on network assumptions, homophily could affect links between any two random nodes, regardless of their spatial position [36]. In a spatial network, however, homophily needs to be redefined concerning space. Therefore, we propose a potential function based on the dissonance value of agents’ opinions, where homophily acts as a gravitational pulling force. Homophilic agents actively move and seek sources of information that optimize this value to a minimal disagreement. This way of modeling homophily is in line with the aforementioned functional purpose of homophily–reducing disagreement [32]. We later argue that this definition relates to the minimum surprise principle. As we showed in our previous work [41], the spatial embedding of information and spatial network is reflected in the collective pattern formation. The result in space is an emerging complex collective behavior, called contour-capturing. Compared to the non-spatial models, echo chambers did not emerge unconditionally in the specific settings of our previous work.
@@
Figure 1: Collective contour capturing emerges as the result of collective opinion dynamics and homophily in space. The traces of agents start with black and become blue (and brighter) over time. Agents converge from the initial random distribution to the mean contour lines of the environment for four different information landscapes. Each plot on the left side corresponds to an information distribution on the right gray scale plot, with the iso-contour lines shown in light yellow and the mean contour line (ground truth) in thick blue line.
1.4 Our Approach: Messengers
@@
To address the problem of echo chambers in our specific scenario, we introduce a new concept of ‘Messengers.’ They are allowed to cross the information space without sampling and serve as data ferries to virtually connect sub-populations across wider distances beyond the communication range limits. In non-spatial networks, a significant parameter in promoting the formation of echo chambers is the limited access of agents to attitude-challenging information [30], i.e. information that deviates from the current belief of the agent. In our model with spatial networks, reducing the communication range would leave the same effect. In this paper, we study how a limited communication range leads to the emergence of echo chambers, in our spatial scenario. Our results confirm that homophily indeed hinders the flow of information, and prevents the collective from achieving a global consensus. To get out of the local optimal trap of echo chambers, agents need to receive diverse opinions flowing into their homogeneous clusters. It might be achieved when agents can communicate with neighbors out of their filter bubble or can escape from the pull of homophily. Thus, as a solution, we define a new state for agents, called Messenger, where they could move freely in space and share their information with others. We investigate if this solution can potentially increase the effective communication range and bring the consensus back to the collective. A challenge, however, is to have a decentralized mechanism for agents to decide whether, when and for how long agents should take a state, and then follow the behavior rules accordingly. Here, we propose and describe a simple decentralized solution based on the Dichotomous Markov Process (DMP) [42]. We study the effect of DMP parameters on individual and collective behaviors, measured by two properties: the ratio of Messengers, and the switching speed between the two states. We then explore the phase diagram of collective behavior and elaborate on an example of collective behavior at each phase. In addition, we highlight a tradeoff between the generalization against specialization of agents and a tradeoff on the population share of Messengers in the collective.
@@ 2 Results
In this section, we investigate the complex behavior of collective consensus resulting from the interplay between opinion dynamics, originating from conformity, and homophily, which is modeled by the movement of agents in space. We then study the emergence of echo chambers in low connectivity regimes and explore a potential solution to restore consensus based on the introduction of an additional state of individuals, which we refer to as Messenger, and a Dichotomous Markov Process (DMP) governing the stochastic transitions between the ordinary state of socially interacting agents and messengers.
2.1 From Consensus to Dissensus in Space: The Emergence of Echo Chambers
As shown previously in [41], the coupling of the local information aggregation and homophily not only allows agents to estimate the mean value of information distribution in the environment but also leads to an emergent spatial collective formation in the physical space (see Fig. 1), referred to as collective contour-capturing behavior. The resulting patterns show that in seeking specific information sources, individuals actively move in space, and collectives form communities based on the information they receive. Here, we explore the result of such complex behavior in more detail and show that global convergence relies on a sufficiently large communication range. In settings with limited connectivity, however, a breakdown to only a local consensus is observed.
@@ 2.1 From Consensus to Dissensus in Space: The Emergence of Echo Chambers
The result of limited connectivity and homophily is that the system becomes trapped into local minima, leading to the formation of echo chambers, with negative consequences on the collective performance due to the inhibition of information flow across the clusters with dissimilar opinions (see Fig. 2).
Additionally, the results demonstrate that echo chambers block the collective movement and prevent it from achieving a consensus in space, i.e. the collective contour-capturing. The clusters emerging from the social interaction model we present in Sec. 4.1 exhibit spatial properties that are otherwise overlooked. An example of such cluster formation is shown in Fig. 3-a, where the echo chambers appear as concentric arcs of various radii, in this case for radial information distribution. In general, the echo chambers take parallel forms which are also parallel to the iso-contours of the information landscape.
Although the original method demonstrated the ability to achieve the contour capturing performance in different environments [41], it is limited to collectives with sufficient connectivity. To show this, we considered the precision of consensus as well as the number of clusters being formed at the end of the experiment. A consensus in space is achieved when agents are precisely on the same contour line. We illustrate these two metrics in Fig. 2, with the number of clusters and spatial precision error ($E_{\text{P}}^{\text{S}}$, for further information, see Sec. 4.4.) The number of clusters and precision error decrease with the communication range after a critical value ($r_{\text{comm}}\approx 0.3$), highlighting a specific regime where the consensus breaks due to the lack of connectivity. This critical value is the minimum required for the initial network to be fully connected. By comparing the final and initial state of the collective in Fig. 2, we argue that the basic method still enhances both the cohesion and precision of the collective. However, in the collectives with very low connectivity, e.g. $r_{\text{comm}}=0.1$, the precision is diminished. We propose and investigate a solution in the next sections to increase effective connectivity by utilizing information-carrying mobile agents, termed ‘Messengers’. The agents switch between the two states following the DMP, which we examine in the following section.
@@ 2.2 Dichotomous Markov Process as a State Switching Mechanism
The solution we propose in this paper is based on the switching mechanism of the DMP, defined by the two parameters $p_{\text{M}}$ and $p_{\text{E}}$. Fig. 4-a illustrates the time history of a single agent’s state. Changing the parameter pair influences two properties of a single agent behavior: the ratio of time each agent spends in either of the two states; and how quickly they switch between the two states. The latter, known as the sojourn time (see Sec 4.3), refers to the time between two consecutive switches. We illustrate the second time spent in the Exploiter and Messenger states in Fig. 4-a by $\tau^{2}_{\text{M}}$ and $\tau^{2}_{\text{E}}$, respectively. At the collective level, changing the two parameters affects the collective average sojourn time ($\tau_{\text{S}}$), and the ratio of Messengers ($m$). These two properties are depicted in the 2D parameter space in Fig. 4-b, c.
We evaluated the Markov process by running multiple independent Monte-Carlo simulations and validated it by comparing the numerical results with the analytical properties in Eq. 7, and  8. The result of analytical and numerical simulations in Fig. 4-b, c suggests validation of the numerical simulation. The switching speed reaches its maximum in the top-right corner of the parameter space ($\log_{10}p_{\text{E}}\approx 0$, $\log_{10}p_{\text{M}}\approx 0$, see Fig. 4-b) and decreases toward the bottom-left corner (small values of $p_{\text{E}}$ and $p_{\text{M}}$). In the top-left corner of Fig. 4-c, the collective is mainly comprised of the Messengers and the ratio decays diagonally to the bottom-right corner. For the sake of comparison, in the rest of the paper, we refer to the setup without Messengers as the baseline setup. In the following sections, we explain how these two properties determine the collective performance. We evaluate the performance of the collective for different DMP parameters in two terms: consensus in the opinion domain, and consensus in space, i.e. contour-capturing behavior.

### v2

@@ Abstract
Collective estimation is a variant of collective decision-making, where agents need to achieve consensus on a continuous quantity in a self-organized fashion via social interactions. A particularly challenging scenario is a fully distributed collective estimation with strongly constrained, dynamical interaction networks, for example, encountered in real physical space. In such cases, agents first need to explore a spatially distributed signal through movement, then reach a consensus while being able to communicate only with nearby neighbors. Collectives face several challenges in achieving precise estimation consensus, particularly due to complex behaviors emerging from the simultaneous evolution of the agents’ opinions and the interaction network. While homophilic networks may facilitate collective estimation in well-connected networks [1], we show that disproportionate interactions with like-minded neighbors lead to the emergence of echo chambers, preventing collective consensus. Our agent-based simulation results confirm that, besides a lack of exposure to attitude-challenging opinions, seeking reaffirming information entraps agents in echo chambers. In a potential solution, agents can break free from the pull of echo chambers. We suggest an additional state where stubborn mobile agents (called Messengers) carry data and connect the disconnected clusters by physically transporting their opinions to other clusters to inform and direct the other agents. However, an agent requires a switching mechanism to determine which state to adopt. We propose a generic, novel approach based on a Dichotomous Markov Process (DMP). We show that a wide range of collective behaviors arise from the DMP. We study a continuum between task specialization with no switching (full-time Messengers), generalization with slow switching (part-time Messengers), and rapid task switching (short-time Messengers). Our results show that stubborn agents can, in various ways, help the collective escape local minima, break the echo chambers, and promote consensus in collective opinion dynamics.
@@ 1 Introduction
Collective behaviors exhibited by animal collectives or groups of interacting artificial agents are fascinating examples of self-organization. Through these behaviors, groups can solve problems collectively, that cannot be solved by any individual alone. This phenomenon is often referred to as collective or swarm intelligence [2, 3, 4]. While significant progress has been made over the past decades in understanding the fundamentals of collective intelligence, many questions remain open regarding actual mechanisms underlying collectively intelligent behavior, in particular in spatially embedded systems lacking the capability for global information exchange between individual agents.
Among the many different manifestations of collective intelligence, the wisdom of crowds effect [5, 6, 7] stands out as a great candidate for studying collective computational intelligence [4]. While the core idea–that the average of many imperfect estimations can be remarkably close to the true value–seems straightforward, achieving precise estimations depends on meeting certain criteria [6, 8, 9, 10, 11]. Typically, it assumes global knowledge by individuals, which is often not achievable in fully decentralized settings. However, distributed consensus models, particularly DeGroot-like models [12], implement the wisdom of crowds effect without a centralization assumption [13, 8]. These models have also been used to model opinion dynamics, by providing a mechanism for how opinions spread across the interaction network [14, 15, 16, 17]. Achieving a consensus has shown potential applications in collective estimation scenarios and can be used in a variety of engineering fields ranging from machine learning to networks of sensors and groups of robots [18, 19, 1]. Moreover, studying these behaviors gives insights into similar complex systems, as used in economics, politics, and social sciences [20, 15, 8, 21, 22]. Individual-based models of opinion dynamics usually follow a simple set of rules; Examples are Voter models, bounded confidence, and gossip methods [23, 24, 25]. The underlying network of interaction among agents is an important determinant of collective behavior [25, 26]. For example, we know that the connectivity of the network controls the tradeoff between speed and accuracy of consensus formation in collectives with static interaction networks–stronger connected networks promote faster consensus at the expense of decision accuracy [1, 11].
@@ 1.1 Homophily and Echo Chambers
Contrary to the assumption of static networks in early models of opinion dynamics, many systems display dynamic networks, as agents tend to constantly rearrange their connections with their peers, making the collective behavior more complex [27]. These rearrangements can be described by specific rules governing the temporal evolution of networks, with homophily being a prominent example. It corresponds to the disproportional tendency of agents to establish links with like-minded neighbors [28, 29, 30]. By adding a new dynamic to the interactions, and hence, altering how the information flows in the network, homophily changes the resulting collective behavior. One potential consequence of homophily in network dynamics is the emergence of ‘echo chambers’: clusters that are highly homogeneous internally and very heterogeneous compared to each other. The strong effects of homophily and echo chambers on collectives highlight their critical importance; particularly, when these effects lead to the emergence of phenomena that are considered threats to human society [31, 32]. An example is their effect on the spread of misinformation [33, 34], where the homogeneity of information within clusters acts as a driver to the diffusion of misinformation [31, 35, 36]. Ultimately, echo chambers reinforce only the locally dominating perspective, foster confirmation biases, and prevent clusters from receiving attitude-challenging information [35, 37, 33]. Other consequences of homophily include the formation of filter bubbles and network segregation [21, 36]. Despite their adverse effects at the collective level, these echo chambers function for a purpose on the individual level: comforting agents with reaffirmation and protecting them from disagreement [35]. Although the actual adversarial outcome of echo chambers in collective dynamics is arguable [38], they indeed inhibit the flow of information on the network in our scenario where achieving consensus is an objective.
1.2 Scenario: Spatial Embedding and Online Sampling of Information
@@
In this work, we consider the scenario of a spatial embedding of opinion dynamics, encountered in the real-world in groups of mobile agents moving through physical space [39, 1]. Many models have already been developed to describe a mechanistic process of how the co-evolution of homophilic networks and opinion dynamics lead to the emergence of echo chambers and complex collective behavior [40, 41, 42, 43]. Although the opinion-updating models coupled to network dynamics vary greatly in the literature, they all typically assume that the initial opinions of agents come from a random pool of possible (discrete) options. Hence, these models do not focus on the origin of the opinions, but rather on their evolution through social interactions. Therefore, by design, these models do not take into account the influence of information aggregation from environmental sources on the evolution of opinions. Contrarily, we assume that the information agents receive is given by a specific information landscape, for example, emitted by one or multiple information sources, and the agent’s position determines the information available to them, hence shaping their opinion. We already know that the position of agents in the network matters [44], but so does the position within the information landscape. The explicit consideration of information space in our model allows us to account for its effects on shaping opinions. Here, the information landscape is modeled by a distribution of environmental signal in 2D Euclidian space, similar to our previous work [1]. Compared to the previous discrete models of initial opinions in homophilic networks, the information landscape provides continuous opinions with arbitrary possible distributions.
@@ 1.3 A Motion Model: Homophily as a Pulling Force
Given our agents are embedded in an information landscape, we require new rules to implement homophily in space. Similarly, we define neighborhood as the local proximity (determined by communication range) of agents in the information landscape. This definition of communication range compares to the thresholds in bounded confidence models of opinion dynamics [24, 45]. In the original models of homophily without spatially constrained networks, homophily could affect links between any two random nodes, regardless of their spatial distance [40]. In a spatial network, however, homophily needs to be redefined to account for spatial locality. Therefore, we propose a potential function based on the dissonance value of agents’ opinions, where homophily acts as a local, gravitational pulling force. Homophilic agents actively move and seek sources of information that optimize this value to a minimal disagreement. This way of modeling homophily is in line with the aforementioned functional purpose of homophily–reducing disagreement [35]. As we showed in our previous work [39], the spatial embedding of information is reflected in the collective pattern formation. In particular, the interplay of individual-level sensing of the information field together with the homophilic interaction enables distributed estimation of the global average of the information landscape and results in a self-organized contour-catching behavior, i.e., aggregation of individuals at the iso-lines of the field value [1] (see Fig.1). However, under certain conditions, such as limited communication range, the spatial network can fragment, leading to the formation of ‘echo chambers.’ In non-spatial networks, a significant parameter in promoting the formation of echo chambers is the limited access of agents to attitude-challenging information [33], i.e., information that deviates from the current belief of the agent. In our model with spatial networks, reducing the communication range has a similar effect. In this paper, we study how a limited communication range leads to the emergence of echo chambers, in our spatial scenario.
@@
Figure 1: Collective contour capturing emerges as the result of collective opinion dynamics and homophily in space. The traces of agents start with black and become blue (and brighter) over time. Agents converge from the initial random distribution to the mean contour lines of the environment for four different information landscapes. Each plot on the left side corresponds to an information distribution on the right grayscale plot, with the iso-contour lines shown in light yellow and the mean contour line (ground truth) in thick blue line. The video of agents optimizing dissonance function following homophily is available via this link: https://figshare.com/s/8c91526b0270c4df8be5
1.4 Breaking-up Echo Chambers with Messengers
@@
To escape the local optimal trap of echo chambers, agents need to receive diverse opinions flowing into their homogeneous clusters. To address the problem in our specific scenario, we introduce a new concept of ‘Messengers’ agents (see details below). These agents serve effectively as data ferries that connect sub-populations across wider distances beyond the communication range limits. It might be achieved when agents can communicate with neighbors out of their filter bubble, or if they can escape from the pull of homophily. Thus, as a solution, besides the ordinary ‘Exploiter’ state, where agents permanently update their opinions by exploiting the information landscape and interacting with others, we introduce a new ‘Messenger’ state. Messenger agents can freely move in space and share their current information with others, but do not acquire any new information, i.e., they do not integrate the opinion of their neighbors and do not sample the information landscape. We investigate if this approach can increase the effective communication range and restore collective consensus. A challenge, however, is to have a decentralized mechanism for agents to decide whether, when, and for how long agents should remain in the Messenger state. Here, we propose a simple decentralized solution based on the Dichotomous Markov Process (DMP) [46]. We study the effect of DMP parameters on individual and collective behaviors, measured by two properties: the ratio of Messengers, and the switching speed between the two states. We then explore the phase diagram of collective behavior and elaborate on an example of collective behavior at each phase.
@@ 2 Results
In this section, first, we review consensus formation in the baseline model with only Exploiter agents and study the emergence of echo chambers in low connectivity regimes. We then investigate the ability to restore consensus via the introduction of an additional state of individuals, which we refer to as Messenger, and a DMP governing the stochastic switching between the Exploiter and the Messenger states. The model details are given in the Methods Sec. 4.1.
2.1 Dissensus in Space: The Emergence of Echo Chambers
@@
As shown previously in [39], the coupling of the local information aggregation and homophily not only allows agents to estimate the mean value of information distribution in the environment but also leads to an emergent spatial collective pattern formation in the information landscape (see Fig. 1.) We refer to this spatial consensus as collective contour-capturing behavior, i.e., the aggregation of agents at the average iso-lines of the information landscape. The resulting patterns show that in seeking specific information sources, individuals actively move in space, and form communities based on the information they receive. Here, we explore the result of such complex behavior in more detail and show that global convergence relies on a sufficiently large communication range. In settings with limited connectivity, however, global consensus breaks down, and only local consensus can be observed. Due to limited connectivity and homophily, the system becomes trapped into local minima, leading to the formation of clusters that serve as echo chambers, with negative consequences on the collective performance due to the inhibition of information flow across the clusters with dissimilar opinions (see Fig. 2).
@@
Our results demonstrate that these clusters inhibit collective movement and prevent the system from achieving consensus in space, i.e., the collective contour-capturing. The clusters form partial contours each aligned locally with the iso-line of the information landscape corresponding to the within-cluster estimation of the average value of the information landscape. An example of such cluster formation is shown in Fig. 3-a, where the echo chambers appear as concentric arcs of various radii, in this case for radial information distribution; while in Fig. 3-b we see a consensus pattern, where agents form a single contour at the average value of the information landscape.
To quantify how the consensus, and the resulting contour-capturing behavior, rely on sufficient connectivity, we measured the number of clusters being formed as well as the precision of consensus ($E_{\text{P}}^{\text{S}}$, for further information, see Sec. 4.2) at the end of the experiment. A consensus in space is achieved when agents are precisely on the same contour line. We illustrate these two metrics in Fig. 2. The number of clusters and precision error decrease with the communication range after a critical value ($r_{\text{comm}}\approx 0.3$), highlighting a specific regime where the consensus breaks due to the lack of connectivity. This critical value is the minimum required for the initial network to be fully connected. By comparing the final and initial state of the collective in Fig. 2, we observe that the basic method still improves both the cohesion and precision of the collective. In the next sections, we propose and investigate a solution for increasing effective connectivity by utilizing information-carrying mobile agents, ‘Messengers’, with the switching in and out of the Messenger state for individual agents being governed by the DMP.
@@ 2.2 Dichotomous Markov Process as a State Switching Mechanism
DMP is a simple stochastic process with the switching dynamics defined by only two parameters $p_{\text{M}}$ and $p_{\text{E}}$. Fig. 4-a illustrates the temporal dynamics of a single agent’s state. Changing the parameter pair influences two properties of a single agent behavior: the ratio of time each agent spends in either of the two states; and how quickly they switch between the two states. The latter, known as the sojourn time (see Methods Sec. 4.1.4), refers to the time between two consecutive switches. We illustrate the time spent in the Exploiter and Messenger states in Fig. 4-a by $\tau_{\text{M}}$ and $\tau_{\text{E}}$, respectively. At the collective level, changing the two parameters affects the collective average sojourn time ($\tau_{\text{S}}$), and the ratio of Messengers ($m$). These two properties are depicted in the 2D parameter space in Fig. 4-b, and Fig. 4-c.
We evaluated the Markov process by running multiple independent Monte-Carlo simulations and validated it by comparing the numerical simulation results to the known analytical solution given in Eq. 7 (also shown in Fig. 4-b, and Fig. 4-c). The switching speed reaches its maximum in the top-right corner of the parameter space, where $p_{\text{M}}$ and $p_{\text{E}}$ are both large ($\log_{10}p_{\text{E}}\approx 0$ and $\log_{10}p_{\text{M}}\approx 0$). It decreases by moving toward the bottom-left corner (small values of $p_{\text{E}}$ and $p_{\text{M}}$). In the top-left corner of Fig. 4-c, the collective is mainly comprised of the Messengers, and the ratio decreases diagonally toward the bottom-right corner. In the following sections, we will explore how the two DMP parameters determine the collective performance in terms of the consensus in the opinion domain, and consensus in information landscape, i.e., contour-capturing behavior. For the rest of the paper, we will refer to the setup without Messengers as the baseline setup.